In the middle of Pacific Ocean is located the small island nation of Tropical Paradise and they use“shells” as their c...

  1. In the middle of Pacific Ocean is located the small island nation of Tropical Paradise and they use“shells” as their currency. The economy of this nation is very small compared to the US. Its autonomous consumption is 400 thousand shells, marginal propensity to consume is 0.9, and total planned investment is 500 thousand shells.

    Calculate their national income using

    1. The Keynesian cross approach

    2. The leakage/injection approach
Answer #1

E 1 Y-AS-45 lei 0 Keynesian cross Approach Y-AL-CAI y = 400+0.94+500 AE =900t 0.94 9000 0.14 = 900 / همه به ooo -۶ 900 9000 &